# VideoPoet: A Large Language Model for Zero-Shot Video Generation

### v3

@@ Abstract
We present VideoPoet, a model for synthesizing high-quality videos from a large variety of conditioning signals. VideoPoet employs a decoder-only transformer architecture that processes multimodal inputs – including images, videos, text, and audio. The training protocol follows that of Large Language Models (LLMs), consisting of two stages: Pretraining and task-specific adaptation. During pretraining, VideoPoet incorporates a mixture of multimodal generative objectives within an autoregressive Transformer framework. The pretrained LLM serves as a foundation that is adapted to a range of video generation tasks. We present results demonstrating the model’s state-of-the-art capabilities in zero-shot video generation, specifically highlighting the generation of high-fidelity motions.
@@ 1 Introduction
Recently, there has been a surge of generative video models capable of a variety of video creation tasks. These include text-to-video (Zhang et al., 2023a; Singer et al., 2022), image-to-video (Yu et al., 2023d), video-to-video stylization (Chen et al., 2023b; Chai et al., 2023; Voleti et al., 2022), and video editing (Ceylan et al., 2023; Wang et al., 2023b; Geyer et al., 2023) among other video applications. Most existing models employ diffusion-based methods for video generation. These video models typically start with a pretrained image model, such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023), that produces high-fidelity images for individual frames, and then fine-tune the model to improve temporal consistency across video frames.
@@
Experiments show VideoPoet’s state-of-the-art capabilities in generating videos with large and high-fidelity motions. Through the powerful capabilities of the transformer architecture, VideoPoet can be straightforwardly trained on a multi-task, multimodal generative objective, allowing for generating consistent and realistic motion driven by text or other prompts. Furthermore, VideoPoet can synthesize coherent long videos of up to 10 seconds by autoregressively extending the content, conditioned on the last second of the generated video.
We also demonstrate that VideoPoet is capable of zero-shot video generation. We use the term “zero-shot video generation” as VideoPoet processes new text, image, or video inputs that diverge from the training data distribution. Furthermore, VideoPoet handles new tasks not included in its training. For example, VideoPoet is able to perform new editing tasks by sequentially chaining training tasks together. The main contributions of this work are:
  • A method for training a Large Language Model (LLM) specifically for video generation tasks, utilizing tokenized video data that incorporates both text-paired and unpaired video data.
  • An approach to video super-resolution that increases spatial resolution within the latent token space using a bidirectional transformer with efficient windowed local attention.
@@ 2 Related Work
As language models can easily incorporate multiple training tasks, task selection is an important area of research. GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2022) demonstrate that training LLMs on diverse tasks leads to positive scaling effects on zero- and few-shot tasks. Other approaches show that masking approaches are a valuable learning target (Hoffmann et al., 2022; Yu et al., 2023a, c). As the model size grows, training data must grow as well (Hoffmann et al., 2022) to maintain similar performance. Our pretraining strategy enables using the same video for multiple training tasks even without paired text. This design facilitates training on a large quantity of video-only examples, thereby decreasing the demand for video-text pairs.
@@ 3.1 Tokenization
We employ the MAGVIT-v2 (Yu et al., 2023c) tokenizer for joint image and video tokenization, and the SoundStream (Zeghidour et al., 2021) tokenizer for audio. Visual and audio vocabularies are concatenated into a unified vocabulary. The text modality is represented by embeddings.
@@
Visual tokenizer is key to generating high-quality video content, often determining the upper limit of achievable video generation quality (Yu et al., 2023c). After analyzing existing tokenizers (Esser et al., 2020; Villegas et al., 2022; Yu et al., 2023a, b), we choose the MAGVIT-v2 (Yu et al., 2023c) tokenizer due to its performance in visual quality and high compression capabilities, which effectively reduce the sequence length required by the LLM, thereby facilitating more efficient and effective learning. Specifically, a video clip is encoded and quantized into a sequence of integers, with a decoder mapping them back to the pixel space. MAGVIT-v2 tokenizes 17-frame 2.125-second 128$\times$128 resolution videos sampled at 8 fps to produce a latent shape of $(5,16,16)$, which is then flattened into 1280 tokens, with a vocabulary size of $2^{18}$. We also tokenize videos into a portrait aspect ratio at 128$\times$224 resolution, producing a latent shape of $(5,28,16)$, or 2240 tokens.
@@
We tokenize audio clips with a pretrained SoundStream (Zeghidour et al., 2021) tokenizer. We embed 2.125 seconds of audio to produce 106 latent frames with a residual vector quantizer (RVQ) of four levels. To improve audio generation performance, we transpose the clip before flattening so that the model predicts the full audio clip at each RVQ granularity level before moving on to the finer grained levels. Finally, each RVQ level has a disjoint vocabulary with each level containing 1,024 codes. This results in a combined audio vocabulary size of 4,096 codes.
@@ 3.3 Super-Resolution
Generating high-resolution (HR) videos with an autoregressive transformer entails heavy computational costs due to the increase in sequence length. To illustrate this with an example, the video tokenizer of Section 3.1 operating on a $17\times 896\times 512$ video produces a sequence of $35,840$ tokens, making autoregressive sampling highly impractical. Aiming at efficient and high-quality generative video upsampling, we develop a custom spatial super-resolution (SR) non-autoregressive video transformer (Yu et al., 2023a) to operate in token space on top of the language model output. To mitigate the computational requirements of the very long sequences involved, and in particular the quadratic memory of the self-attention layers, our design incorporates windowed local attention (Gupta et al., 2022). Specifically, our SR transformer is composed of blocks of three transformer layers, each of which performs self-attention in a local window aligned with one of three axes (Tu et al., 2022): spatial vertical, spatial horizontal and temporal. The cross-attention layers attend to the low-resolution (LR) token sequence and are also divided into local windows, isomorphic to those of the self-attention layers. All blocks also include cross-attention to T5 XL text embeddings. See Fig. 3 for a schematic representation of the custom transformer architecture.
Similar to (Yu et al., 2023c), we train the SR transformer using token factorization (with $k=2$ factors) to account for the large vocabulary size. The LR token sequences are obtained by tokenizing bicubic-downsampled versions of the ground truth videos and applying noise augmentation (Ho et al., 2022a) in the discrete latent space. Specifically, we randomly resample the value of a random subset of the LR tokens and independently drop the LR condition and text embeddings for 10% of the training samples. During inference, we use non-autoregressive sampling (Chang et al., 2022; Yu et al., 2023a) with classifier-free guidance independently on both the LR condition and the text embeddings (Brooks et al., 2023). We use a cascade of two 2×2$\times$2 × stages to generate videos of $896\times 512$ resolution from the $224\times 128$ base output of VideoPoet. More implementaiton details can be found in the appendix.
@@ 4.1 Task Prompt Design
We consider the following tasks. Unconditioned video generation: Generate video frames without conditioning on an input. Text-to-video (T2V): Generate video from a text prompt. Video future prediction (FP): Given an input video of variable length, predict future frames. Image-to-video (I2V): Given the first frame of a video as an input image, predict the future frames. Video inpainting/outpainting (Painting): Given a masked video, predict the video with the masked contents filled in. Video stylization: Given text, optical flow, and depth, predict the video frames (Section 4.1). Audio-to-video: Given an input audio waveform, predict the corresponding video. Video-to-audio: Given an input video, predict the corresponding audio waveform. Audio-video continuation (AVCont) given an input frame and its audio, predict the rest of the video and audio.
@@ 4.2 Training Strategy
We fine-tune our pretrained model for enhanced performance on specific tasks or for new task adaptation, such as text-to-video and image-to-video tasks, using a high-quality data subset. This results in improved generation quality, consistent with Zhou et al. (2023), and addresses decoding collapse issues, characterized by repetitive token predictions. Such fine-tuning not only diversifies outputs but also allows for a higher classifier-free guidance scale (Ho & Salimans, 2022), boosting overall quality.
@@ 5.1 Experimental Setup
We adopt widely used metrics such as Fréchet Video Distance (FVD) (Unterthiner et al., 2018), CLIP similarity score (Wu et al., 2021), and Inception Score (IS) (Saito et al., 2020) for evaluation. Note that the specific metrics and evaluation methods vary across different datasets. Detailed information on these variations can be found in Section A.5.4. We include examples of the generated videos in the supplementary materials.
@@ 5.2 Pretraining Task Analysis
The evaluation results are presented in Table 1. We assess a model across the four tasks within the zero-shot evaluation benchmark: the T2V task on MSR-VTT (Xu et al., 2016) and UCF 101 (Soomro et al., 2012), the FP on K600 (Carreira et al., 2018), and central inpainting and outpainting on SSv2 (Goyal et al., 2017). In these experiments, we employ a single model to perform all the tasks. The model is not trained on the training data of these evaluation datasets, and thus it is a zero-shot evaluation.
The top rows of Table 1 depict each pretraining task configuration of the 300 million parameter model, which are comparable in their setups. Our evaluation benchmarks span diverse visual domains, posing a challenge to achieving consistent improvement across all of them. Nevertheless, incorporating all pretraining tasks results in the best overall performance, on average, across all evaluated tasks. Additionally, the significant disparity observed in the “SSL” row suggests the limitations of self-supervised training and underscores the necessity for text-paired data during training. The last row, “ALL (8B)”, is the model with 8 billion parameters, trained on the pretraining tasks as discussed in Section 3 and utilized significantly more compute.
@@ 5.3 Comparison with the State-of-the-Art
We analyze VideoPoet using human raters and compare with other recent models: Show-1 (Zhang et al., 2023a), VideoCrafter (Chen et al., 2023a), Phenaki (Villegas et al., 2022), Pika (Pika, 2023), Gen2 (Runway, 2023) and Lumiere (Bar-Tal et al., 2024). Show-1, VideoCrafter, Pika, Gen2 and Lumiere are video diffusion models while Phenaki is a token-based model using masked token modeling (Chang et al., 2022). We ran the most up-to-date model versions as of January 2024.
@@
For this user study we use the fine-tuned version of VideoPoet as discussed in Section 4.2 and compare against alternative models in a side-by-side fashion for each prompt. Raters are shown videos generated by two models at a time (in randomized order so as to not bias raters). Not all methods generate videos at the same size or aspect ratio, and we resize each video to a fixed area while maintaining its original aspect ratio. Raters are then asked to compare the videos along 5 dimensions and for each dimension to report which video is better. The 5 dimensions are: (1) text fidelity (which video follows the text prompt most faithfully), (2) video quality, (3) motion “interestingness”, (4) motion realism, and (5) temporal consistency. Raters are required to go through a collection of training examples for each of these 5 dimensions before they begin.
Our findings are summarized in Fig. 4, where green and pink bars represent the proportion of trials where VideoPoet was preferred or less preferred over an alternative, respectively. We observe that VideoPoet outperforms all baseline models along almost all of the dimensions. More specifically, VideoPoet achieves significant wins along the motion categories (motion interestingness and realism, temporal consistency) and Lumiere (Bar-Tal et al., 2024) which is diffusion based and concurrent to our work, is the only model that outperforms VideoPoet on Video Quality.
@@ A.2 Model Scale and Performance
[…]
@@ A.2.1 Qualitative Comparison of 1B and 8B models
In Figure A.2.1, we show outputs of 1B and 8B parameter models on the same prompts. Four frames from the best video output of each model in a batch of four text-to-video samples were selected to represent the model. In the first row, the 1B model is unstable with large changes to the subject over time and misses elements from the complex prompt. This prompt was originally used for scaling comparisons in (Yu et al., 2022), and compared to a dedicated image-only model, our model does not preserve text as well given the training data used. In the second row, we use a simpler text task and show that the 8B model can represent a single letter clearly, but the 1B model still produces artifacts. In the third row, we show that the 8B model learns spatial positioning such as the river being in front of the astronaut and horse. In the fourth row, we show that the 8B parameter model learned a stop motion style to have items disappear “one by one” and can follow a complicated layout from a long prompt. In contrast, the 1B model includes all of the nouns, but is unstable over time and does not follow the layout indicated in the prompt. In the bottom row, we show that the 8B model understands counts of objects in that it displays a full bouquet (though 12 roses are not explicitly in frame) and smooth consistent motion as opposed to the 5 roses and distorting objects produced by the 1B model. Overall, scaling the model improved temporal consistency, prompt fidelity, and motion dynamics while adding capabilities for limited text rendering, spatial understanding, and counting.

### v4

@@ Abstract
We present VideoPoet, a model for synthesizing high-quality videos from a large variety of conditioning signals. VideoPoet employs a decoder-only transformer architecture that processes multimodal inputs – including images, videos, text, and audio. The training protocol follows that of Large Language Models (LLMs), consisting of two stages: pretraining and task-specific adaptation. During pretraining, VideoPoet incorporates a mixture of multimodal generative objectives within an autoregressive Transformer framework. The pretrained LLM serves as a foundation that is adapted to a range of video generation tasks. We present results demonstrating the model’s state-of-the-art capabilities in zero-shot video generation, specifically highlighting the generation of high-fidelity motions. Project page: https://sites.research.google/videopoet/.
@@ 1 Introduction
Recently, there has been a surge of generative video models capable of a variety of video creation tasks. These include text-to-video (Zhang et al., 2023a; Singer et al., 2022), image-to-video (Yu et al., 2023c), video-to-video stylization (Chen et al., 2023b; Chai et al., 2023; Voleti et al., 2022), and video editing (Ceylan et al., 2023; Wang et al., 2023b; Geyer et al., 2023) among other video applications. Most existing models employ diffusion-based methods for video generation. These video models typically start with a pretrained image model, such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023), that produces high-fidelity images for individual frames, and then fine-tune the model to improve temporal consistency across video frames.
@@
Experiments show VideoPoet’s state-of-the-art capabilities in generating videos with large and high-fidelity motions. With the powerful capabilities of the transformer architecture, VideoPoet can be straightforwardly trained on a multi-task, multimodal generative objective, allowing for generating consistent and realistic motion driven by text or other prompts. Furthermore, VideoPoet can synthesize coherent long videos of up to 10 seconds by autoregressively extending the content, conditioned on the last second of the generated video.
We also demonstrate that VideoPoet is capable of zero-shot video generation. We use the term “zero-shot video generation” as VideoPoet processes new text, image, or video inputs that diverge from the training data distribution. Furthermore, VideoPoet handles new tasks not included in its training. For example, VideoPoet is able to perform new editing tasks by sequentially chaining tasks together.
Our contribution is a proof of concept demonstrating an understudied approach to high-quality video generation with LLMs, distinct from the dominant diffusion-based methods. Specifically, the main contributions include:
  • A method for training a Large Language Model (LLM) specifically for video generation, utilizing tokenized data that incorporates both text-paired and unpaired videos.
  • A video super-resolution method that increases spatial resolution within the latent token space using a bidirectional transformer with efficient windowed local attention.
@@ 2 Related Work
As language models can easily incorporate multiple training tasks, task selection is an important area of research. GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2022) demonstrate that training LLMs on diverse tasks leads to positive scaling effects on zero- and few-shot tasks. Other approaches show that masking approaches are a valuable learning target (Hoffmann et al., 2022; Yu et al., 2023a, 2024). As the model size grows, training data must grow as well (Hoffmann et al., 2022) to maintain similar performance. Our pretraining strategy enables using the same video for multiple training tasks even without paired text. This design facilitates training on a large quantity of video-only examples, thereby decreasing the demand for video-text pairs.
@@ 3.1 Tokenization
We employ the MAGVIT-v2 (Yu et al., 2024) tokenizer for joint image and video tokenization, and the SoundStream (Zeghidour et al., 2021) tokenizer for audio. Visual and audio vocabularies are concatenated into a unified vocabulary. The text modality is represented by embeddings.
@@
Visual tokenizers are key to generating high-quality video content, often determining the upper limit of achievable video generation quality (Yu et al., 2024). Among existing tokenizers (Esser et al., 2020; Villegas et al., 2022; Yu et al., 2023a, b), we choose the MAGVIT-v2 (Yu et al., 2024) tokenizer due to its performance in visual quality and high compression capabilities, which effectively reduce the sequence length required by the LLM, thereby facilitating more efficient and effective learning. Specifically, a video clip is encoded and quantized into an integer sequence integers, with a decoder mapping back to pixel space. MAGVIT-v2 tokenizes 17-frame 2.125-second 128$\times$128 resolution videos sampled at 8 fps to produce a latent shape of $(5,16,16)$, which is then flattened into 1280 tokens, with a vocabulary size of $2^{18}$. We also tokenize videos into portrait aspect ratio at 128$\times$224 resolution, producing a latent shape of $(5,28,16)$, or 2240 tokens.
@@
We tokenize audio clips with a pretrained SoundStream (Zeghidour et al., 2021) tokenizer. We embed 2.125 seconds of audio to produce 106 latent frames with a residual vector quantizer (RVQ) of four levels. Two possible choices exist for predicting the audio tokens with the RVQ representation: 1) sequentially predicting the entire audio clip from a lower to a higher RVQ level, or 2) simultaneously predicting all RVQ levels for a single audio token. Our results suggest that the former method demonstrates a slight advantage over the latter approach. Finally, each RVQ level has a disjoint vocabulary with each level containing 1,024 codes. This results in a combined audio vocabulary size of 4,096 codes.
@@ 3.3 Super-Resolution
Generating high-resolution (HR) videos autoregressively entails heavy computational costs due to the increase in sequence length. To illustrate this with an example, the video tokenizer of Section 3.1 operating on a $17\times 896\times 512$ video produces a sequence of $35,840$ tokens, making autoregressive sampling highly impractical. Aiming at efficient and high-quality generative video upsampling, we develop a custom spatial super-resolution (SR) non-autoregressive video transformer (Yu et al., 2023a) to operate in token space on top of the language model output. To mitigate the computational requirements of the very long sequences involved, and in particular the quadratic memory of the self-attention layers, our design incorporates windowed local attention (Gupta et al., 2022). Specifically, our SR transformer is composed of blocks of three transformer layers, each of which performs self-attention in a local window aligned with one of three axes (Tu et al., 2022): spatial vertical, spatial horizontal and temporal. The cross-attention layers attend to the low-resolution (LR) token sequence and are also divided into local windows, isomorphic to those of the self-attention layers. All blocks also include cross-attention to T5 XL text embeddings. See Fig. 3 for a schematic representation of the custom transformer architecture.
To account for the larger vocabulary size, we follow (Yu et al., 2024), and train the SR transformer using token factorization with $k=2$ factors, which converts a $262,144$-way classification problem into two $512$-way classification problems. The LR token sequences are obtained by tokenizing bicubic-downsampled versions of the ground truth videos and applying noise augmentation (Ho et al., 2022a) in the discrete latent space, to mitigate the distribution mismatch between real and generated videos. Specifically, we randomly resample the value of a random subset of the LR tokens and independently drop the LR condition and text embeddings for 10% of the training samples. During inference, we use non-autoregressive sampling (Chang et al., 2022; Yu et al., 2023a) with classifier-free guidance independently on both the LR condition and the text embeddings (Brooks et al., 2023). We use a cascade of two 2×2$\times$2 × stages to generate videos of $896\times 512$ resolution from the $224\times 128$ base output of VideoPoet. More implementaiton details can be found in the appendix.
@@ 4.1 Task Prompt Design
We consider the following tasks. Unconditioned video generation: Generate video frames without conditioning on an input. Text-to-video (T2V): Generate video from a text prompt. Video future prediction (FP): Given an input video of variable length, predict future frames. Image-to-video (I2V): Given the first frame of a video as an input image, predict the future frames. Video inpainting/outpainting (Painting): Given a masked video, predict the video with the masked contents filled in. Video stylization: Given text, optical flow, and depth, predict the video frames (Section 4.1). Audio-to-video: Given an input audio waveform, predict the corresponding video. Video-to-audio: Given an input video, predict the corresponding audio waveform. Audio-video continuation (AVCont) given an input frame and its audio, predict the rest of the video and audio. In principle, the model can generate text, but we have not explicitly evaluated this ability.
@@ 4.2 Training Strategy
We fine-tune our pretrained model for enhanced performance on specific tasks or for new task adaptation, such as text-to-video and image-to-video tasks, using a data subset of higher quality. These videos are sourced from broad internal sources with millions of videos that contain simpler clips, and are not manually tailored or selected. This results in improved generation quality, consistent with Zhou et al. (2023), and addresses decoding collapse issues, characterized by repetitive token predictions. Such fine-tuning not only diversifies outputs but also allows for a higher classifier-free guidance scale (Ho & Salimans, 2022), boosting overall quality.
@@ 5.1 Experimental Setup
We employ widely used metrics such as Fréchet Video Distance (FVD) (Unterthiner et al., 2018), CLIP similarity score (Wu et al., 2021), and Inception Score (IS) (Saito et al., 2020) for evaluation. Note that the specific metrics and evaluation methods vary across different datasets. Detailed information on these variations can be found in Section A.5.5. We include examples of the generated videos in the supplementary materials.
@@ 5.2 Pretraining Task Analysis
The evaluation results are presented in Table 1. We assess a model across the four tasks within the zero-shot evaluation benchmark: the T2V task on MSR-VTT (Xu et al., 2016) and UCF 101 (Soomro et al., 2012), the FP on K600 (Carreira et al., 2018), and central inpainting and outpainting on SSv2 (Goyal et al., 2017). The audio generation results (FAD) are reported in Fig. 5(b) of the Appendix. In these experiments, we employ a single model to perform all the tasks. The model is not trained on the training data of these evaluation datasets, and thus it is a zero-shot evaluation.
The top rows of Table 1 depict each pretraining task configuration of the 300 million parameter model, which are comparable in their setups. Our evaluation benchmarks span diverse visual domains, posing a challenge to achieving consistent improvement across all of them. Nevertheless, incorporating all pretraining tasks results in the best overall performance, on average, across all evaluated tasks. Additionally, the significant disparity observed in the “SSL” row suggests the limitations of self-supervised training and underscores the necessity for text-paired data during training. Both single-task and multi-task models are trained for the same number of steps. The minor decrease in performance of multi-task training in Table 1 might be due to the insufficient training of each task. The last row, “ALL (8B)”, is the model with 8 billion parameters, trained on the pretraining tasks as discussed in Section 3 and utilized significantly more compute.
@@ 5.3 Comparison with the State-of-the-Art
We analyze VideoPoet using human raters and compare with other recent models: Show-1 (Zhang et al., 2023a), VideoCrafter (Chen et al., 2023a), Phenaki (Villegas et al., 2022), Pika (Pika, 2023), Gen2 (Runway, 2023), WALT (Gupta et al., 2023) and Lumiere (Bar-Tal et al., 2024). Show-1, VideoCrafter, Pika, Gen2, WALT and Lumiere are video diffusion models while Phenaki is a token-based model using masked token modeling (Chang et al., 2022). We ran the most up-to-date model versions as of January 2024 and note that WALT and Lumiere were concurrently developed and not available during initial submission of this paper.
@@
For this user study we use the fine-tuned version of VideoPoet as discussed in Section 4.2. As part of our system we also use a fixed negative prompt for all inference calls as well as lightweight prompt rewriting. We then compare VideoPoet against alternative models in side-by-side fashion for each prompt. Raters are shown videos generated by two models at a time (in randomized order so as to not bias raters). We refer readers to Appendix A.5.4 for more details of our methodology.
Not all methods generate videos at the same size, aspect ratio, or framerate, and we thus resize and resample each video to a fixed area while maintaining its original aspect ratio as well as common framerate. Raters are then asked to compare the videos along 5 dimensions and for each dimension to report which video is better. The 5 dimensions are: (1) text fidelity (which video follows the text prompt most faithfully), (2) video quality, (3) motion “interestingness”, (4) motion realism, and (5) temporal consistency. Raters are required to go through a collection of training examples for each of these 5 dimensions before they begin.
Our findings are summarized in Fig. 4, where green and pink bars represent the proportion of trials where VideoPoet was preferred or less preferred over an alternative, respectively. We observe that VideoPoet is competitive with state of the art video diffusion models despite taking a dramatically different approach, even outperforming these baseline models along most dimensions. More specifically, VideoPoet achieves significant wins along motion interestingness and realism and Lumiere (Bar-Tal et al., 2024) which is diffusion based and concurrent to our work, is the only model that outperforms VideoPoet on Video Quality.
5.4 Runtime
For a batch size of 4 videos and generating 17 frames at 8fps using TPUv5p (4 chips) accelerators, our base model runs in 34s, the detokenizer (converting tokens to pixels) requires 1.3s and super-resolution is 6.8s — thus, amortized run time is about 5 seconds per second of output video. Note that our model has not been optimized, and any acceleration techniques applicable to standard LLMs could be applied here as well.
@@ Acknowledgements
Impact Statement
This paper introduces research aimed at advancing the field of video generation and introduces a new tool to enhance human creativity. In considering the societal impact of our video generation model, VideoPoet, several concerns emerge, including the potential for misuse and ethical considerations. Misuse may entail the creation of deceptive or harmful content, such as misinformation or deepfakes. Ethical considerations encompass ensuring that generated content adheres to ethical standards, avoids perpetuating harmful stereotypes, and respects cultural sensitivities. To mitigate these risks, we adhere to evolving strategies developed within our community. This includes using digital watermarking in generated videos to enable traceability and accountability, and maintain transparency in our model design to foster trust.
@@ A.2 Model Scale and Performance
(a) Video generation quality in FVD ($\downarrow$).
(b) Audio generation quality in FAD ($\downarrow$).
@@ A.2.1 Qualitative Comparison of 1B and 8B models
In Figure 5, we show outputs of 1B and 8B parameter models on the same prompts. Four frames from the best video output of each model in a batch of four text-to-video samples were selected to represent the model. In the first row, the 1B model is unstable with large changes to the subject over time and misses elements from the complex prompt. This prompt was originally used for scaling comparisons in (Yu et al., 2022), and compared to a dedicated image-only model, our model does not preserve text as well given the training data used. In the second row, we use a simpler text task and show that the 8B model can represent a single letter clearly, but the 1B model still produces artifacts. In the third row, we show that the 8B model learns spatial positioning such as the river being in front of the astronaut and horse. In the fourth row, we show that the 8B parameter model learned a stop motion style to have items disappear “one by one” and can follow a complicated layout from a long prompt. In contrast, the 1B model includes all of the nouns, but is unstable over time and does not follow the layout indicated in the prompt. In the bottom row, we show that the 8B model understands counts of objects in that it displays a full bouquet (though 12 roses are not explicitly in frame) and smooth consistent motion as opposed to the 5 roses and distorting objects produced by the 1B model. Overall, scaling the model improved temporal consistency, prompt fidelity, and motion dynamics while adding capabilities for limited text rendering, spatial understanding, and counting.